# HRNet: Differentially Private Hierarchical and Multi-Resolution Network for Human Mobility Data Synthesization

### v1

@@ 3.1. Overview
We provide the overview of multi-task training of HRNet, illustrated in Figure 2. HRNet builds upon the baseline model that models conditional probabilities. See Section 2.3 for detail of the baseline model. Note that we simplify our explanation, sidelining the training of time information which is the same as the baseline, to emphasize the principal distinctions.
Consider a scenario depicted in Figure 2. The example involves two resolutions, featuring $4$ and $16$ grid cells respectively, and trajectory $(1,2,6,10)$ at resolution $2$. In the learning phase for $v=(1,2,6,10)$, particularly at step $3$ (step $1$ is learning $2$ from prefix ($1$), step $2$ is learning $6$ from prefix $(1,2)$, and step $3$ is learning $10$ from prefix $(1,2,6)$), HRNet operates as follows:
@@ 3.2.1. The Structure
The core of HRNet’s location encoding is a trainable root vector, denoted as $\theta_{\rm root}$. This root vector is augmented with a 2D quad deconvolutional network comprising of $d$ layers. The process involves the application of this 2D quad deconvolutional network to $\theta_{\rm root}$, producing $4^{d}$ distinct vectors. Essentially, this operation results in an embedding matrix of dimensions $(2^{d},2^{d},n_{\rm dim})$:
$M_{d}={\rm deconv}^{d}(\theta_{\rm root}).$
For a detailed explanation of the ${\rm deconv}$ operation, refer to Section 2.5. Notably, the number of parameters in each deconvolutional layer only depends on the kernel size and is independent of $n_{\rm POI}$, sustaining a constant number of parameters. Consequently, this architecture requires only $O(d)=O(\log n_{\rm POI})$ parameters to generate $n_{\rm POI}$ vectors, which is a significant reduction compared to the $O(n_{\rm POI})$ parameters typically needed in embedding matrix.
@@ 3.2.2. Vector Assignment
The assignment of the $4^{d}$ vectors generated by the hierarchical location encoding component to the $n_{\rm POI}$ POIs follows a specific methodology. Given that the grid size of a POI, $w*w$, is a power of 4, we set $d=\log_{4}(w*w)$. This approach facilitates an alignment of vectors with the grid, described as:
@@
Here, we assumed that POIs are aligned according to a uniform grid as described in Section 2.2.1 for simplicity. Note that this grid assumption is not necessary as long as the number of scattered POIs is equal to or less than $4^{d}$ since the assignment method is arbitrary. However, as mentioned above, we should follow the first law of geography rather than randomly assigning vectors to the scattered POIs. Thus, we propose a method for assigning scattered POIs to vectors in $M_{d}$ following the proximity principle. First, we create a grid that covers all scattered POIs. Here, we set the minimum $d$ so that the number of POIs is less than $4^{d}$, and $w=2^{d}$. Then, we compute the coordinates of the center of each grid cell (${\rm cell}_{j}$) and assign each POI (${\rm POI}_{i}$) to a cell so that the Euclidean distance $d({\rm POI}_{i},{\rm cell}_{j})$ is as small as possible. This assignment can be formalized as a linear assignment problem with a cost matrix $C$, where
@@
This optimization can be solved using the Jonker-Volgenant algorithm (Crouse, 2016). Once each POI is assigned to a grid cell, we can use the vector assignment method described above. Since POIs are aligned based on geographical proximity, this embedding also incorporates the proximity principle of the deconvolutional network.
@@ 3.3. Scoring
In this process, the encoded prefix sequence $(v_{1},\dots,v_{i-1})$, represented by $h_{i-1}$, is transformed into a query vector. This transformation is executed using a feed-forward neural network, denoted as $f_{\rm query}$. Concurrently, the encoded POI, ${\rm hiencode}(l)$, is processed into a key vector through another feed-forward neural network, named $f_{\rm key}$. Mathematically, this can be represented as ${\rm query}=f_{\rm query}(h_{i})$ and ${\rm key}_{l}=f_{\rm key}({\rm hiencode}(l)).$ Following these transformations, the score of $l$ as $v_{i}$ is determined by computing the dot product between the query and the key vector ${\rm score}_{l}={\rm query}\cdot{\rm key}_{l}.$ The final step involves applying the softmax function to these scores across all POIs, resulting in a probability distribution over POIs:
$\Pr(v_{i}=l_{i}|v_{1},\dots,v_{i-1})={\rm softmax}({\rm score}_{L})[l_{i}],$
where ${\rm score}_{L}=({\rm score}_{1},\dots,{\rm score}_{|L|})$.
@@ 3.4.1. Multi-resolution Interpretation
We consider grid divisions of $2^{i_{\text{res}}}\times 2^{i_{\text{res}}}$ at each resolution level $i_{\text{res}}\in\{1,2,\dots,d-1\}$. Here, the $i_{\text{res}}$th resolution includes $L_{i_{\text{res}}}$ grid cells, represented as:
$i_{\rm res}\in[d],L_{i_{\rm res}}:=[4^{i_{\rm res}}].$
The matrix $M_{i_{\text{res}}}$ (refer to Section 3.2.1) encodes these grid cells by applying the vector assignment in Section 3.2.2. The encoding for grid cell $l$ at resolution $i_{\text{res}}$ is thus:
@@ 3.4.2. The Objective Function
where ${\rm up}_{i_{\rm res}}(y)\in L_{i_{\rm res}}$ is the grid cell covering $y$ at resolution $i_{\text{res}}$. $\Pr(l_{i_{\rm res}}=l|v_{1},\dots,v_{i-1})$ represents the inferred probability of the next grid cell being $l$ at resolution $i_{\rm res}$. These losses for all resolutions are then summed up, and this sum is used to compute the gradient for backward.
@@ 4.1. DP Transition Matrix
where $\mathbf{1}$ is an indicator function and $V_{l_{i_{\rm res}},l}$ is a subset of trajectory universe $\mathcal{V}$, which includes trajectories that have transition from $l_{i_{\rm res}}$ to $l$:
$V_{l_{i_{\rm res}},l}:=\{\mathbf{v}\in\mathcal{V}|\exists i,{\rm up}_{i_{\rm res}}(\mathbf{v}[i])=l_{i_{\rm res}}\ \&\ \mathbf{v}[i+1]=l\}.$
Note that this calculation normalizes the count with the sequence length $|\mathbf{v}|$, to limit sensitivity to $1$ for DP. To ensure DP, we introduce Laplace noise into the matrix as follows:
@@
where ${\rm Laplace}(\varepsilon)$ is a Laplace noise that leads to $\varepsilon$-DP when sensitivity is $1$. This approach results in a denser transition matrix, resulting in a more informative and less noisy transition matrix that is useful in pre-training. Figure 1 left side shows a sample of such coarse transition matrix.
@@ 4.2. Pre-training
This section outlines the pre-training methodology using the DP transition matrix. During this phase, the core objective is to generate a probability distribution depicting the likelihood of transitions from a given input grid cell to the POIs. It is important to note that while both the hierarchical location encoding component and its scoring component undergo pre-training, the prefix encoding component (GRU) (see Figure 2) does not.
@@ 4.2.1. The Architecture
where $x$ is the input vector that represents the initial grid cell, with further details provided in the following section. Notably, this temporary component is exclusive to the pre-training phase and is replaced with GRU during model training.
@@ 4.2.2. The Objective Function
We propose a model that utilizes an augmentation approach, achieved through the proportional blending of grid cells at resolution $i_{\rm res}$. A mixing ratio vector is defined as: $\mathbf{r}=(r_{1},\dots,r_{|L_{i_{\rm res}}|})\in[0,1]^{|L_{i_{\rm res}}|},$ subject to the constraint $\sum_{i=1}^{|L_{i_{\rm res}}|}r_{i}=1$. Here, each $r_{i}$ denotes the proportion of grid cell $i$ in the initial state, forming a prior distribution as a multinomial distribution with probabilities $\mathbf{r}$. Then, the ground truth probability distribution given $\mathbf{r}$ is calculated as:
@@
where ${\rm Mult}(\cdot|\mathbf{r})$ is the multinomial distribution with probabilities $(r_{1},\dots,r_{|L_{i_{\rm res}}|})$ and $\Pr(l_{\rm next}|l_{\rm initial}=l)={\rm DPTRAN}_{i_{\rm res}}[l]$. This essentially models the probability distribution for the next POI, given a stay at $l_{\rm initial}$ with the probability $\Pr(l_{\rm initial}|\mathbf{r})$. Since $\mathbf{r}\in[0,1]^{|L_{i_{\rm res}}|}$, the ground truths are not confined to a limited number. The mixing ratio is generated by sampling from a Dirichlet distribution $\mathbf{r}\sim{\rm Dirichlet}(\mathbf{a}),$ where $\mathbf{a}=(1,1,\dots,1)$ and $|\mathbf{a}|=|L_{i_{\rm res}}|$.
@@
This is the weighted sum of the encoding of grid cells $l\in L_{i_{\rm res}}$ with ratio $\mathbf{r}$. Then, we use the mixed encoding $x$ as input to the temporal component (i.e., Equation (5)) to derive the key vector and then the output distribution. Finally, the loss is
@@
where $\hat{y}|\mathbf{r}$ is the output distribution derived from the key vector. With this objective function, our model not only learns the probability distributions of ${\rm DPTRAN}$ but also gains a continuous representation among these probability distributions.
4.3. Discussion on The Choice
We proposed pre-training HRNet using DPTRAN (Equation 4) based on the KL loss in Equation (6). Here, we explain the rationale behind this choice from two perspectives:
@@ 4.3. Discussion on The Choice
One challenge of DP-SGD is that it requires many iterations due to starting from a cold state. A warm start helps to mitigate this issue (Amid et al., 2022), and we anticipate this benefit in our approach. The main training phase includes optimizing DPTRAN with the KL loss through multi-task learning. It is important to note that optimizing cross-entropy loss is essentially the same as optimizing KL loss. Although marginalization to generate DPTRAN means the KL loss is not exactly the original cross-entropy loss, we expect it to be similar enough to provide a warm start.
@@
There are several options for DPTRAN beyond the first-order transition, such as second-order transitions and transitions that include time information. Additionally, we could use a sophisticated and ununiform resolution based on density, similar to privtree (Zhang et al., 2016), for the initial grids of DPTRAN. As shown by Zhou et al. (Zhou et al., 2021), using public data to find the gradient subspace addresses the issue of a large number of parameters and improves noise stability. Therefore, a more precise DPTRAN as mentioned above could enhance the main training results. However, these methods typically require additional privacy budget. From our experience, we found that using first-order transition and a fixed resolution ($i_{\rm res}=2$) is more beneficial than these complex methods. First-order transitions are denser than second-order transitions and those with time information, making the perturbed information more useful for pre-training. A fixed resolution allows us to allocate a larger portion of the privacy budget to perturb the transition matrix, which provides better information for pre-training. Moreover, a fixed resolution enables algorithmic privacy budget allocation (see Section 4.4.1) without seeing raw data.
Remark
@@ 4.4. Privacy Analysis of HRNet with Pre-training
Here, ${\rm dptran}(D,\varepsilon_{2})$ refers to the computation of Equation (4) and pretrain is the GD training with loss function (6). DPSGD is the algorithm described in Section 2.1, utilizing the allocated privacy parameters $(\varepsilon_{1},\delta)$, and initial parameter $\theta^{\prime}$. Deriving the trained parameters for HRNet (i.e., $\theta^{\prime\prime}$) satisfies $(\varepsilon_{1}+\varepsilon_{2},\delta)$-DP.
@@
We can see that the entire computation is the sequential composition of Equation (7) and Equation (8). For Equation (7) (i.e., the pre-training phase), the raw data $D$ is employed solely for generating DPTRAN (i.e., Equation (4)). The application of the Laplace mechanism, which is known to satisfy $\varepsilon_{2}$-DP (Dwork et al., 2006), ensures that the pre-training phase is $\varepsilon_{2}$-DP. According to the post-processing theorem (Dwork et al., 2014), computing $\theta^{\prime}$ satisfies $\varepsilon_{2}$-DP since it does not further utilize $D$. For Equation (8) (i.e., the training phase), the raw data $D$ is used for DPSGD and DPSGD satisfies $(\varepsilon_{1},\delta)$-DP (see Section 2.1 for detail). By applying the composition theorem of DP (Dwork et al., 2014), deriving the trained parameters for HRNet (i.e., $\theta^{\prime\prime}$) satisfies $(\varepsilon_{1}+\varepsilon_{2},\delta)$-DP. ∎
@@ 4.4.1. Privacy Budget Allocation
Assume that we use $i_{\rm res}$th resolution for the initial grid cell of the DP transition matrix. We propose the following allocation formula:

### v2

@@ 3.1. Overview
We provide the overview of multi-task training with hierarchical location encoding of HRNet, illustrated in Figure 2. HRNet builds upon the baseline model that models conditional probabilities. See Section 2.3 for detail of the baseline model. Note that we simplify our explanation, sidelining the training of time information which is the same as the baseline, to emphasize the principal distinctions.
Consider a scenario depicted in Figure 2. The example involves two resolutions, featuring $4$ and $16$ grid cells respectively, and trajectory $(1,2,6,10)$ at resolution $2$. In the learning phase for $v=(1,2,6,10)$, step $1$ is learning $2$ from prefix ($1$), step $2$ is learning $6$ from prefix $(1,2)$, and step $3$ is learning $10$ from prefix $(1,2,6)$. Consider step 3, HRNet operates as follows:
@@ 3.2.1. The structure
The core of HRNet’s location encoding is a trainable root vector, denoted as $\theta_{\rm root}$. This root vector is augmented with a 2D quad deconvolutional network comprising of $d$ layers. The process involves the application of this 2D quad deconvolutional network to $\theta_{\rm root}$, producing $4^{d}$ distinct vectors. Essentially, this operation results in an embedding matrix of dimensions $(2^{d},2^{d},n_{\rm dim})$: $M_{d}={\rm deconv}^{d}(\theta_{\rm root}).$ For a detailed explanation of the ${\rm deconv}$ operation, refer to Section 2.5. Notably, the number of parameters in each deconvolutional layer only depends on the kernel size and is independent of $n_{\rm POI}$, sustaining a constant number of parameters. Consequently, this architecture requires only $O(d)=O(\log n_{\rm POI})$ parameters to generate $n_{\rm POI}$ vectors, which is a significant reduction compared to the $O(n_{\rm POI})$ parameters typically needed in embedding matrix.
@@ 3.2.2. Vector assignment
The assignment of the $4^{d}$ vectors generated by the hierarchical location encoding component to the $n_{\rm POI}$ POIs follows a specific methodology. Given that the grid size, $w*w$, is a power of 4, we set $d=\log_{4}(w*w)$. This approach facilitates an alignment of vectors with the grid, described as:
@@
The above assignment assumed that POIs are the grid cells according to a uniform grid as described in Section 2.2.1 for simplicity. Note that this grid assumption is not necessary and HRNet can work with scattered POIs as long as the number of POIs is equal to or less than $4^{d}$. As mentioned above, we should follow the first law of geography rather than randomly assigning vectors to the scattered POIs. Thus, we propose a method for assigning scattered POIs to vectors in $M_{d}$ following the proximity principle. First, we create a grid that covers all scattered POIs. Here, we set the minimum $d$ so that the number of POIs is less than $4^{d}$, and $w=2^{d}$. Then, we compute the coordinates of the center of each grid cell (${\rm cell}_{j}$) and assign each POI (${\rm POI}_{i}$) to a unique cell so that the Euclidean distance $d({\rm POI}_{i},{\rm cell}_{j})$ is as small as possible. This assignment can be formalized as a linear assignment problem with a cost matrix $C$, where
@@
This optimization can be solved using the Jonker-Volgenant algorithm (Crouse, 2016). Once each POI is assigned to a unique grid cell, we can use the vector assignment method described above. Since POIs are aligned based on geographical proximity, this embedding also incorporates the proximity principle of the deconvolutional network.
@@ 3.3. Scoring
In this process, the encoded prefix sequence $(v_{1},\dots,v_{i-1})$, represented by $h_{i-1}$, is transformed into a query vector. This transformation is executed using a feed-forward neural network, denoted as $f_{\rm query}$. Concurrently, the encoded POI, ${\rm hiencode}(l)$, is processed into a key vector through another feed-forward neural network, named $f_{\rm key}$. Mathematically, this can be represented as ${\rm query}=f_{\rm query}(h_{i})$ and ${\rm key}_{l}=f_{\rm key}({\rm hiencode}(l)).$ Following these transformations, the score of $l$ as $v_{i}$ is determined by computing the dot product between the query and the key vector ${\rm score}_{l}={\rm query}\cdot{\rm key}_{l}.$ The final step involves applying the softmax function to these scores across all POIs, resulting in a probability distribution over POIs: $\Pr(v_{i}=l_{i}|v_{1},\dots,v_{i-1})={\rm softmax}({\rm score}_{L})[l_{i}],$ where ${\rm score}_{L}=({\rm score}_{1},\dots,{\rm score}_{|L|})$.
@@ 3.4.1. Multi-resolution Interpretation
We consider grid divisions of $2^{i_{\text{res}}}\times 2^{i_{\text{res}}}$ at each resolution level $i_{\text{res}}\in\{1,2,\dots,d-1\}$. Here, the $i_{\text{res}}$th resolution includes $L_{i_{\text{res}}}$ grid cells, represented as: $i_{\rm res}\in[d],L_{i_{\rm res}}:=[4^{i_{\rm res}}].$ The matrix $M_{i_{\text{res}}}$ (refer to Section 3.2.1) encodes these grid cells by applying the vector assignment in Section 3.2.2. The encoding for grid cell $l$ at resolution $i_{\text{res}}$ is thus:
@@ 3.4.2. The Objective Function
where ${\rm up}_{i_{\rm res}}(y)\in L_{i_{\rm res}}$ is the grid cell covering $y$ at resolution $i_{\text{res}}$. $\Pr(l_{i_{\rm res}}=l|v_{1},\dots,v_{i-1})$ represents the inferred probability of the next grid cell being $l$ at resolution $i_{\rm res}$. These losses for all resolutions are then summed up, and this sum is used to compute the gradient for backward propagation.
@@ 4.1. DP Transition Matrix
where $\mathbf{1}$ is an indicator function and $V_{l_{i_{\rm res}},l}$ is a subset of trajectory universe $\mathcal{V}$, which includes trajectories that have transition from $l_{i_{\rm res}}$ to $l$: $V_{l_{i_{\rm res}},l}:=\{\mathbf{v}\in\mathcal{V}|\exists i,{\rm up}_{i_{\rm res}}(\mathbf{v}[i])=l_{i_{\rm res}}\ \&\ \mathbf{v}[i+1]=l\}.$ Note that this calculation normalizes the count with the sequence length $|\mathbf{v}|$, to limit sensitivity to $1$ for DP. To ensure DP, we introduce Laplace noise into the matrix as follows:
@@
where ${\rm Laplace}(\varepsilon)$ is a Laplace noise that leads to $\varepsilon$-DP when sensitivity is $1$. This approach results in a denser transition matrix, hence more informative and less noisy, which is useful in pre-training. Figure 1 left side shows a sample of such coarse transition matrix.
@@ 4.2. Pre-training
This section outlines the pre-training methodology using the DP transition matrix. During this phase, the core objective is to generate a probability distribution depicting the likelihood of transitions from a given region (a grid cell at resolution $i_{\rm res}$) to the POIs. It is important to note that while both the hierarchical location encoding component and its scoring component undergo pre-training, the prefix encoding component (GRU) (see Figure 2) does not since the DP transition matrix does not contain prefix information.
@@ 4.2.1. The Architecture
where $x$ is the input vector that represents the given region, with further details provided in the following section. Notably, this temporary component is exclusive to the pre-training phase and is replaced with GRU during model training.
@@ 4.2.2. The Objective Function
To enhance learning, we propose a data augmentation approach to augment the transition matrix, achieved through proportional blending of regions (grid cells at resolution $i_{\rm res}$). A mixing ratio vector is defined as: $\mathbf{r}=(r_{1},\dots,r_{|L_{i_{\rm res}}|})\in[0,1]^{|L_{i_{\rm res}}|},$ subject to the constraint $\sum_{i=1}^{|L_{i_{\rm res}}|}r_{i}=1$. Here, each $r_{i}$ denotes the proportion of region $i$ in the initial state, forming a prior distribution as a multinomial distribution with probabilities $\mathbf{r}$. Then, the ground truth probability distribution given $\mathbf{r}$ is calculated as:
@@
where ${\rm Mult}(\cdot|\mathbf{r})$ is the multinomial distribution with probabilities $(r_{1},\dots,r_{|L_{i_{\rm res}}|})$ and $\Pr(l_{\rm next}|l_{\rm initial}=l)={\rm DPTRAN}_{i_{\rm res}}[l]$. This essentially models the probability distribution for the next POI, given a stay at $l_{\rm initial}$ with the probability $\Pr(l_{\rm initial}|\mathbf{r})$. The mixing ratio is generated by sampling from a Dirichlet distribution $\mathbf{r}\sim{\rm Dirichlet}(\mathbf{a}),$ where $\mathbf{a}=(1,1,\dots,1)$ and $|\mathbf{a}|=|L_{i_{\rm res}}|$. Note that the augmented transition matrix is still DP due to the post processing property.
@@
This is the weighted sum of the encoding of region $l\in L_{i_{\rm res}}$ with ratio $\mathbf{r}$. Then, we use the mixed encoding $x$ as input to the temporal component (i.e., Equation (5)) to derive the key vector and then the output distribution. Finally, the loss is
@@
where $\hat{y}|\mathbf{r}$ is the output distribution derived from the key vector. With this objective function, our model not only learns the probability distributions of ${\rm DPTRAN}$ but also learns from a continuous representation among these probability distributions.
4.2.3. Discussion
In summary, HRNet is pretrained using DPTRAN (Equation 4) based on the KL loss (Equation 6). Here, we further explain the rationale behind this choice from two perspectives:
@@ 4.2.3. Discussion
One challenge of DP-SGD is that it requires many iterations due to starting from a cold state. A warm start helps to mitigate this issue (Amid et al., 2022), and we anticipate this benefit in our approach. This is because when the KL loss during pre-training is optimized, it also helps to optimize the cross-entropy loss during main training due to the multi-task learning. It is important to note that optimizing cross-entropy loss is essentially the same as optimizing KL loss. Although marginalization to generate DPTRAN means the KL loss is not exactly the original cross-entropy loss, we expect it to be similar enough to provide a warm start.
@@
There are several alternatives for pretraining beyond the first-order transition matrix DPTRAN, such as second-order transitions and transitions that include time information. Additionally, we could use a sophisticated and nonuniform grid based on density, similar to privtree (Zhang et al., 2016), for the initial grid of DPTRAN. As shown by Zhou et al. (Zhou et al., 2021), using public data to find the gradient subspace addresses the issue of a large number of parameters and improves noise stability. Therefore, a more precise DPTRAN as mentioned above could potentially enhance the main training results. However, these methods typically require additional privacy budget. From our experience, we found that using first-order transition and a fixed resolution ($i_{\rm res}=2$) is more beneficial than these complex methods. First-order transitions are denser than second-order transitions and those with time information, making the perturbed information more useful for pre-training. A fixed resolution allows us to allocate a larger portion of the privacy budget to perturb the transition matrix, which provides better information for pre-training. Moreover, a fixed resolution enables algorithmic privacy budget allocation (see Section 4.3.1) without seeing raw data.
@@ 4.3. Privacy Analysis of HRNet with Pre-training
Here, ${\rm dptran}(D,\varepsilon_{2})$ refers to the computation of Equation (4) and pretrain is the standard training with loss function (6). DPSGD is the private training algorithm described in Section 2.1, utilizing the allocated privacy parameters $(\varepsilon_{1},\delta)$, and initial parameter $\theta^{\prime}$. Deriving the trained parameters for HRNet (i.e., $\theta^{\prime\prime}$) satisfies $(\varepsilon_{1}+\varepsilon_{2},\delta)$-DP.
@@
We can see that the entire computation is the sequential composition of Equation (7) and Equation (8). For Equation (7) (i.e., the pre-training phase), the raw data $D$ is employed solely for generating DPTRAN (i.e., Equation (4)). The application of the Laplace mechanism, which is known to satisfy $\varepsilon_{2}$-DP (Dwork et al., 2006), ensures that the pre-training phase is $\varepsilon_{2}$-DP. According to the post-processing theorem (Dwork et al., 2014), computing $\theta^{\prime}$ satisfies $\varepsilon_{2}$-DP since it does not further utilize $D$. For Equation (8) (i.e., the training phase), the raw data $D$ is used for DPSGD and DPSGD satisfies $(\varepsilon_{1},\delta)$-DP (see Section 2.1 for detail). By applying the composition theorem of DP (Dwork et al., 2014), deriving $\theta^{\prime\prime}$ satisfies $(\varepsilon_{1}+\varepsilon_{2},\delta)$-DP. ∎
@@ 4.3.1. Privacy Budget Allocation
Assume that we use the $i_{\rm res}$th resolution for the initial region of the DP transition matrix. We propose the following allocation formula: